# Online Item Cold-Start Recommendation with Popularity-Aware Meta-Learning

### v2

@@ Abstract.
With the rise of e-commerce and short videos, online recommender systems that can capture users’ interests and update new items in real-time play an increasingly important role. In both online and offline recommendation, the cold-start problem due to interaction sparsity has been affecting the recommendation effect of cold-start items, which is also known as the long-tail problem of item distribution. Many cold-start scheme based on fine-tuning or knowledge transferring shows excellent performance on offline recommendation. Yet, these schemes are infeasible for online recommendation on streaming data pipelines due to different training method, computational overhead and time constraints. Inspired by the above questions, we propose a model-agnostic recommendation algorithm called Popularity-Aware Meta-learning (PAM), to address the item cold-start problem under streaming data settings. PAM divides the incoming data into different meta-learning tasks by predefined item popularity thresholds. The model can distinguish and reweight behavior-related and content-related features in each task based on their different roles in different popularity levels, thus adapting to recommendations for cold-start samples. These task-fixing design significantly reduces additional computation and storage costs compared to offline methods. Furthermore, PAM also introduced data augmentation and an additional self-supervised loss specifically designed for low-popularity tasks, leveraging insights from high-popularity samples. This approach effectively mitigates the issue of inadequate supervision due to the scarcity of cold-start samples. Experimental results across multiple public datasets demonstrate the superiority of our approach over other baseline methods in addressing cold-start challenges in online streaming data scenarios.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXX.XXXXXconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining ; August 3-7, 2025; Toronto, Canadaisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Platforms that depends on online recommendation service, including short video, e-commerce, and streaming media, play an indispensable role in people’s lives. Online recommender systems need to capture interest shifts in the system through periodic updates while taking into account the historical information of users and items. Many previous works (Devooght et al., 2015; Zhang et al., 2020; Wang et al., 2020) have demonstrated superior performance in the realm of online recommendation. These online systems are characterized by a large amount of data, strict time-consuming constraints, changes in global information (e.g. user interests, item popularity) over time (Peng et al., 2021) and the requirement of one epoch streaming training. A common way to train the system streamingly is continuously optimizing the model to capture information shifts. The same network parameters are applied for all users and items, and the parameters are updated using data arriving during the period to avoid being stale.
@@
Such cold-start problem has garnered considerable attention and has been addressed through various solutions in recommendation scenarios. The representation-based approach (Lee et al., 2019; Kang et al., 2019; Zhu et al., 2021) considers items as tasks and generates personalized parameters for different items to enhance the effectiveness of cold-start item recommendation, such as meta-learning (Lee et al., 2019) and knowledge transfer (Zhu et al., 2021; Chen et al., 2023). The side-information-based approach (Pan et al., 2022; Yu et al., 2018; Chen et al., 2024) focuses on using content-based side information and pay less attention to the item’s popularity-related information, which also have different importance in the recommendation of cold and popular items. However, the methods above are still challenging to apply in online scenarios with streaming data.
The cold-start problem in online recommendations faces two main challenges. On the one hand, in the combination of online and cold-start recommendation, the Matthew effect (Merton, 1968) makes the system optimize the distribution of popular items further. Because the popular items make up the majority of the training data leading to better performance on popular items, thus the data further increases the proportion of popular data over time. This effect exacerbates the problem of cold-starts in online systems. Meanwhile, the requirement of training streaming data with one epoch makes the existing cold-start methods inapplicable, and the time consumption and computational demand make generating personalized parameters for items impractical in online scenarios.
@@ 4.3.2. Data Augmentation
With relatively low traffic proportion, the cold-start task in meta-learning still has limited training opportunity. Data argumentation can significant increase the number of cold-start samples, so we propose a simulation-based data argumentation. We treat the simulated cold-start embedding $\hat{\boldsymbol{e}}_{i}$ from the previous subsection as a new part of cold-start task data with the real popular interaction data labels $y_{ui}$, and train the meta-learning model to update the cold-start task parameters, which can partially augment the data to improve the effect.
@@ 4.3.3. Self-supervised Instructor
Algorithm 1 The training and test process of PAM.
@@
  Randomly initialize parameters $\Phi_{0}$, $\Theta_{0}$, $\alpha_{0}$, $f^{Sup}_{0}$
@@
     $\mathcal{L}^{M}_{t}=\sum_{n=1}^{k}\lambda_{n}\mathcal{L}_{T_{n}}(\Omega_{t}^{n},\Phi_{t}^{n}|{D_{t}^{n}}^{Q})$triangleright Main meta-learning loss
@@ 5. Experiments
We conducted experiments to answer the following RQs: RQ1: How does PAM’s performance on online cold-start recommendations improve over existing frameworks? RQ2: How much of a performance enhancement do the various components of PAM provide? RQ3: Why PAM perform better on multitasking? RQ4: How do the various hyperparameters of PAM affect the performance results? RQ5: How does PAM perform in online A/B testing?
@@
Dataset #Users #Items #Inter. #Tags #Sparsity
@@ 5.1.1. Datasets
We selected three public datasets that containing timestamp to ensure that the characteristic of online streaming data is simulated, including MovieLens (Harper and Konstan, 2015), Yelp and Book (Hou et al., 2024). Details of datasets are shown in  A.
@@ 5.1.2. Baselines
- SML (Zhang et al., 2020) proposes a meta-model that takes the previous moment’s recommender and current moment’s data as inputs and generate the next moment’s recommender for serving through a CNN-based network.
@@ 5.6. Online A/B Tests (RQ5)
To evaluate the performance of PAM in an online system, we deployed the PAM in a commercial, billion-user-scale, online recommender system and compared it to the PF baseline. For company privacy, we don’t report the implementation details and the real performance of the original retrieval stage. Instead, we report the performance gain ratio improved by our approach PAM. In the online system, we counted the following metrics, including the rate of items appearing in users’ recommendations (Show%), and the ratio of users liking (LTR), commenting (CMTR), and collecting videos (CLTR). The results of the experiments in Table 4 show that the high accuracy of PAM for item recommendation enables items to be better recommended to users who favor the item.
@@ Acknowledgements.
This work is partially sponsored by National Key R&D Program of China under Grant 2022YFB3104200, NSFC 62XXXXXX and 62032003.

### v3

@@ Abstract.
With the rise of e-commerce and short videos, online recommender systems that can capture users’ interests and update new items in real-time play an increasingly important role. In both online and offline recommendation systems, the cold-start problem caused by interaction sparsity has been impacting the effectiveness of recommendations for cold-start items. Many cold-start scheme based on fine-tuning or knowledge transferring shows excellent performance on offline recommendation. Yet, these schemes are infeasible for online recommendation on streaming data pipelines due to different training method, computational overhead and time constraints. Inspired by the above questions, we propose a model-agnostic recommendation algorithm called Popularity-Aware Meta-learning (PAM), to address the item cold-start problem under streaming data settings. PAM divides the incoming data into different meta-learning tasks by predefined item popularity thresholds. The model can distinguish and reweight behavior-related and content-related features in each task based on their different roles in different popularity levels, thus adapting to recommendations for cold-start samples. These task-fixing design significantly reduces additional computation and storage costs compared to offline methods. Furthermore, PAM also introduced data augmentation and an additional self-supervised loss specifically designed for low-popularity tasks, leveraging insights from high-popularity samples. This approach effectively mitigates the issue of inadequate supervision due to the scarcity of cold-start samples. Experimental results across multiple public datasets demonstrate the superiority of our approach over other baseline methods in addressing cold-start challenges in online streaming data scenarios.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709336isbn: 979-8-4007-1245-6/25/08ccs: Information systems Recommender systems
@@ 1. Introduction
Platforms that depends on online recommendation service, including short video, e-commerce, and streaming media, play an indispensable role in people’s lives. Online recommender systems need to capture interest shifts in the system through periodic updates while taking into account the historical information of users and items. Many previous works (Devooght et al., 2015; Zhang et al., 2020; Wang et al., 2020) have demonstrated superior performance in the realm of online recommendation. These online systems are characterized by a large amount of data, strict time-consuming constraints, changes in global information over time (Peng et al., 2021) and the requirement of one epoch streaming training. A common way to train the system streamingly is continuously optimizing the model to capture information shifts. The network parameters are applied for all users and items, and the parameters are updated using data arriving in the period to avoid being stale.
@@
Such cold-start problem has garnered considerable attention and has been addressed through various solutions in recommendation scenarios. The representation-based approach (Lee et al., 2019; Kang et al., 2019; Zhu et al., 2021) considers items as tasks and generates personalized parameters for different items to enhance the effectiveness of cold-start item recommendation, such as meta-learning (Lee et al., 2019) and knowledge transfer (Zhu et al., 2021; Chen et al., 2023; Song et al., 2024b, a). The side-information-based approach (Pan et al., 2022; Yu et al., 2018; Chen et al., 2024) focuses on using content-based side information and pay less attention to the item’s popularity-related information, which also have different importance in the recommendation of cold and popular items. However, the methods above are still challenging to apply in online scenarios with the characteristic of streaming data.
The cold-start problem in online recommendation systems presents two main challenges. Firstly, when integrating online and cold-start recommendations, the Matthew effect causes the system to over-recommend popular items. This happens because popular items dominate the training data, leading to improved performance for these items and an increasing proportion of popular data over time. This effect exacerbates the cold-start problem in online systems. Secondly, the influx of numerous cold-start items meaks existing methods ineffective, as the time and computational demands make it impractical to generate personalized parameters for each cold-start item in online scenarios.
@@ 4.3.2. Data Augmentation
With relatively low traffic proportion, the cold-start task in meta-learning still has limited training opportunity. Data augmentation can significant increase the number of cold-start samples, so we propose a simulation-based data augmentation. We treat the simulated cold-start embedding $\hat{\boldsymbol{e}}_{i}$ from the previous subsection as a new part of cold-start task data with the real popular interaction data labels $y_{ui}$, and train the meta-learning model to update the cold-start task parameters, which can partially augment the data to improve the effect.
@@ 4.3.3. Self-supervised Instructor
Algorithm 1 The training and serving process of PAM.
@@
  Randomly initialize parameters $\Phi_{0}$, $\Theta_{0}$, $f^{Sup}_{0}$
@@
     $\mathcal{L}^{M}_{t}\leftarrow\sum_{n=1}^{k}\lambda_{n}\mathcal{L}_{T_{n}}(\Omega_{t}^{n},\Phi_{t}^{n}|{D_{t}^{n}}^{Q})$triangleright Meta-learning loss
@@ 5. Experiments
We conducted experiments  [^1] Our code is available at https://github.com/Sycamoretail/PAM. to answer the following RQs: RQ1: How does PAM’s performance on online cold-start recommendations improve over existing frameworks? RQ2: How much of a performance enhancement do the various components of PAM provide? RQ3: Why PAM perform better on multitasking? RQ4: How do the various hyperparameters of PAM affect the performance results? RQ5: How does PAM perform in online A/B testing?
@@
Dataset #Users #Items #Inter. #Tags Density
@@ 5.1.1. Datasets
We selected three public datasets that containing timestamp to ensure that the characteristic of online streaming data is simulated, including MovieLens (Harper and Konstan, 2015), Yelp and Book (Hou et al., 2024). Details of datasets are shown in appendix A.
@@ 5.1.2. Baselines
- SML (Zhang et al., 2020) proposes a meta-model that takes the previous moment’s recommender and current moment’s data as inputs and generate the next moment’s recommender for serving.
@@ 5.6. Online A/B Tests (RQ5)
To evaluate the performance of PAM in an online system, we deployed the PAM in a commercial, billion-user-scale, online recommender system and compared it to the PF baseline. For company privacy, we don’t report the implementation details such as the size of online data, the QPS of user requests and the real performance of the original retrieval stage. Instead, we report the performance gain ratio improved by our approach PAM. In the online system, we counted the following metrics, including the rate of items appearing in users’ recommendations (Show%), and the ratio of users liking (LTR), commenting (CMTR), and collecting videos (CLTR). The results of the experiments in Table 4 show that the high accuracy of PAM for item recommendation enables items to be better recommended to users who favor the item.
@@ Acknowledgements.
This work is partially sponsored by National Key R&D Program of China under Grant 2022YFB3104200, NSFC U24A20235 and 62032003.